OpenStudy (anonymous):

\[cos^{-1}\left[ cos \frac{\pi}5 \right]=\] Where do I start? I know that \[0 \le \theta \le\frac{\pi}2\] do I find the value of \[ cos \frac{\pi}5\] first?

geerky42 (geerky42):

Since \(\cos^{-1}\) and \(\cos\) is inverse of each other, they cancel each other.

OpenStudy (anonymous):

ooops \[0 \le \theta \le\pi\]

OpenStudy (anonymous):

what about the \[\frac\pi 5\]?

OpenStudy (anonymous):

can I ignore that?

geerky42 (geerky42):

\(\cos\) and \(\cos^{-1}\) both cancel each other, which leave us just \(\dfrac{\pi}{5}\).

geerky42 (geerky42):

Does this help?

OpenStudy (anonymous):

yes. Thank you. It would be the same thing if sin and inverse sine were switched around correct?

OpenStudy (anonymous):

\[sin(sin^{-1}x)\]

geerky42 (geerky42):

Yes.

OpenStudy (anonymous):

as long as it's within the limits I guess